OpenStudy (anonymous):

can anyone subtract x/x^2-1 - 2/x+1

OpenStudy (anonymous):

(-x+2)/(x^2-1)

OpenStudy (anonymous):

To start with you factor (x^2-1) and you get (x+1) (x-1) so it's now x/(x+1)(x-1) - 2/(x+1) in order to subtract these two terms they need a common denominator. so you multiply the second expression by (x-1) to make it look like the first, 2 (x -1) = 2x -2 Now you can subtract both terms, x - 2x +2 = -x +2 write the new term above the denominator, (-x+2)/(x^2-1)

OpenStudy (anonymous):

thank you that helps alot